OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x to the 4th power over x to the second power. Explain in your own words how you can simplify x to the 4th power over x to the second power without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

medals for everyone

OpenStudy (mertsj):

\[\frac{(x)(x)(x)(x)}{(x)(x)}=(x)(x)=x^2\]

OpenStudy (mertsj):

you can simplify without writing the expanded form by subtracting the exponents

OpenStudy (mertsj):

\[\frac{16x^3y^6}{8xy^2}=2x^2y^4\]